# Asymmetric dynamical localization and precision measurement of BEC micromotion

### v1

@@ Abstract
We show that a Bose-Einstein Condensate (BEC) launched with non-zero initial momentum into a periodically kicked optical lattice creates an asymmetrically localized momentum distribution in a moving frame with a small initial current. This asymmetric localization is investigated under two scenarios; (a) when the BEC is in motion in the laboratory frame and, (b) when the optical lattice is in motion in the laboratory frame. The asymmetric features are shown to arise from the early-time dynamics induced by the broken parity symmetry and, asymptotically, freeze as the dynamical localization stabilizes. The micromotion of BEC is measured using the early-time asymmetry. In this context, micromotion refers to the extremely low initial velocity of the BEC along the lattice direction. This originates from the jitter when the hybrid trap potential is turned off. By employing BEC in a kicked and moving optical lattice, the asymmetry in early-time dynamics is measured to precisely characterize and quantify the micromotion phenomena in the quantum system. Micromotion measurement has applications in quantifying systematic shifts and uncertainties in light-pulse interferometers.
@@ I Introduction
In recent years, there has been much interest in studying quantum kicked rotor (QKR) with BEC – with and without tunable interactions [1, 2, 3, 4]. Further, the ease of imprinting the lattice phase on the BEC also provides different types of effects such as on-resonance quantum ratchets [5, 6] and quantum boomerang effect [7, 8, 9]. BEC based QKR has also been utilized in coupled quantum kicked rotors [1] where two incommensurate optical lattices drive a quantum to classical transition by breaking dynamical localization (DL). A discrete-time quantum walk has also been observed in the BEC-based system [10]. Here, we employ a QKR test-bed based on a BEC to experimentally demonstrate the asymmetric dynamical localization and we utilize this feature to perform precision measurement of BEC micromotion.
@@
Precision measurements with light-pulse atom interferometers have opened tremendous applications in quantum sensing [11, 12, 13]. Atom interferometers are successfully utilized in gravimetry [14, 15, 13], rotation sensing [16, 17], magnetometers [18], and the determination of photon recoil [19]. Particularly in precision rotation sensing and gravimetry, nullifying the systematic shifts and measurement errors coming from Coriolis effect [20] in phase shifts necessitates precise knowledge of the initial velocity [16, 21, 14]. This initial velocity may arise from external launch velocities or micromotion coming from small movement of the atomic cloud during the turnoff of the trapping potential. In a rotation sensor, the phase shift caused due to this initial velocity adds to the Sagnac phase shift [20]. The Sagnac phase shift is in the range of tens of milliradians (mrad) for a velocity of 100 $\mu$m/s over an interferogram time of hundreds of milliseconds (ms). The velocity range of micromotion can lie in the range of 100-1000 $\mu$m/s, which is much smaller compared to the velocity imparted due to two-photon recoil momentum ($\approx$ 1 cm/s). The measurement of such a small velocity is challenging and requires huge time-of-flight in standard absorption imaging or very precise Bragg or Raman spectroscopy to measure the Doppler shift [21, 22, 23].
@@
Firstly, we experimentally demonstrate localization in momemtum space by launching the BEC with varying initial momentum or by inducing the lattice motion. A constant phase evolution of the launched wave function or the lattice motion creates an asymmetric momentum distribution. Furthermore, we illustrate that the asymmetric nature of the momentum distribution can be easily controlled by adjusting the launch momentum rather than altering the direction of the launch momentum. In contrast to previous studies [24] with cold atoms, where such asymmetry arises from pulse shape effects and observed after long time, here we demonstrate that such asymmetry arises much earlier in time [7, 5] and the asymmetry gets settled as dynamical localization takes place. An asymmetry in the localized state has been reported in Ref.[7] in a BEC by single phase change and phase reversal in the beginning. By utilizing this asymmetric behaviour of the localization in the early-time dynamics, we present a method to measure the micromotion of the BEC [25, 26, 27] in an Atom Optics Kicked Rotor (AOKR). The AOKR is already utilized for measurement of gravity through survival resonances [28, 29]. We demonstrate that even this micromotion of the BEC can induce early-time asymmetry in the momentum distribution. This asymmetry directly quantifies the velocity of micromotion, aligning with the primary focus of the current study.
II QKR in moving optical lattice
@@
The quantum kicked rotor (QKR) is a fundamental model of quantum chaos [30] extensively explored for its demonstration of dynamical localization [31, 32]. The classical kicked rotor can display chaotic dynamics accompanied by a diffusive growth of mean energy. Over extended time scales, quantum interference effects inhibit classical diffusive dynamics, a phenomenon in momentum space analogous to Anderson localization [33]. Since the first realization of the QKR using cold atoms [34], QKR has spurred numerous experimental investigations to explore a variety of scenarios that inhibit localization [35, 36, 37]. Physically, the standard kicked rotor describes a particle subjected to periodic kicks imparted by the stationary optical lattice created by counter-propagating laser beams. However, in this work, we consider a kicked rotor system in which the optical lattice moves at a constant velocity in the laboratory frame. The Hamiltonian of QKR in a moving lattice is given by [38]
$H=\frac{\hat{p}^{2}}{2}+K\cos(k\hat{x}-\alpha t)~{}\sum_{n=1}^{N}~{}\delta(t-nT),$ (1)
@@
where, $\hat{p}$ and $\hat{x}$ represent the momentum and position operators respectively. They obey canonical commutation relation $[\hat{x},\hat{p}]=i\hbar_{\rm eff}$, where the effective Planck constant $\hbar_{\rm eff}$ can be tuned in the experiment. Further, $K$ is the stochastic parameter, $T$ is the time period between consecutive kicks, $k$ is the wave vector, and $\alpha$ is the frequency difference between two lattice beams. The lattice velocity $v$ arising from the frequency difference between the counter-propagating beams is $v=\lambda\alpha$, where $\lambda$ is the wavelength of the optical lattice. Throughout this work, parameters will be chosen so that the classical analogue of QKR displays chaos, with $K\geq 5$. This parameter choice ensures that the localization effects we observe are of quantum origin.
@@
$U=\exp{\left(-i\frac{p^{2}T}{2}\right)}\exp{\left(-iK\cos(kx-\alpha t)\right)}.$ (2)
@@
and is consistent with the initial distribution of BEC. Here, $\sigma_{w}$ characterizes the width of the wavefunction in position space, while $p_{0}$ is the initial velocity, typically arising from launched velocity or micromotion. Generally, in our experiments, when BEC is launched with initial velocity $p_{0}$, the lattice velocity is stationary and vice versa. In general, either moving the lattice appropriately or the atomic cloud imparted with initial velocity $p_{0}$ in the lab frame are expected to induce similar effects [39, 5].
For numerical simulations, we use the standard split-operator method to evolve an initial state of the kicked rotor. This method consists of two primary components: the kick operator, which is diagonal in position space, and the free evolution operator, which is diagonal in momentum space. To simulate scenarios where the lattice is in motion in the lab frame, we imprint the phase just before each kick. Similarly, to model the situation in which BEC moves in the laboratory frame, we initialize the a wavefunction that incorporates the motion right from the outset.
@@
If the lattice is moved with constant velocity, then the kick potential (in the lab frame) $V(x)=K\cos(kx-\alpha t)$ induces a path difference of $\alpha T$ at each kick. Consequently, the corresponding phase difference between successive kicks is $\phi=2\pi\alpha T$ and this breaks the parity symmetry. The total phase difference accumulated after $n$ kicks is $\phi_{n}=2\pi\alpha T\times(n-1)$, assuming that the phase is initialized to zero for the first kick. The accumulated phase difference $\phi_{n}$ over short timescales induces an inhomogeneity in position space due to broken parity symmetry, leading to an asymmetric momentum distribution [5]. This asymmetry can be quantified through $\langle p(t=2T)\rangle$ immediately following $n=2$ kicks.
In the experiments, we control the lattice velocity by tuning $\alpha$ and subsequently measure $\langle p(t=2T)\rangle$. Based on physical consideration and since phases are unique only upto $2\pi$, we posit that the average momentum after two kicks to have a form
$\langle p(t=2T)\rangle=c~{}\sin(2\pi\alpha T),$ (4)
where $c$ is a constant. This implies that if $\alpha T=n/2$ (where $n$ is an integer), $\langle p\rangle=0$ implying an absence of asymmetry for specific choice of initial velocity and kick period. In the long time limit of $n\gg 1$, the initial asymmetry accumulated in the short-time limit eventually freezes due to the emergence of dynamical localization. Hence, the early-time dynamics dictates the long-term behaviour and the onset of asymmetrical dynamical localization in the system.
For numerical simulations, we use the standard split-operator method to evolve an initial state of the kicked rotor. This method consists of two components: the kick operator (diagonal in position space), and the free evolution operator (diagonal in momentum space). To simulate scenarios where the lattice is in motion in the lab frame, the phase just before each kick is imprinted on the evolving wavefunction.
To gauge micromotion accurately, the optical lattice is precisely moved with frequency difference of the order of $100$Hz, aligning it with the scale of micromotion. Upon achieving a velocity for the lattice that corresponds precisely to the micromotion, we observe $\langle p(t=2T)\rangle$ equal to zero. This alignment establishes a direct correspondence between the velocity of the lattice and the velocity of micromotion.
@@ III Experimental setup for QKR
The QKR setup we have used for this work is similar to the one described here [37]. However, instead of cold atoms, we utilize a BEC of 87Rb every 8 seconds, through forced evaporative cooling. The atoms are initially prepared in the |$\ket{F=1,m_{F}=-1}$ state, with a BEC temperature of $80$ nK and a population of approximately $40,000$ atoms. This BEC serves as the initial wave function for our experimental investigations. The optical standing wave is produced using two independent laser beams, which are generated by a single laser passing through two separate acousto-optic modulators. We can control the frequency difference between lattice beams and their switching. Switching time and laser power provide us control over scaled Planck constant $\hbar_{\rm eff}$ and stochastic parameter $K$.
This work comprises two main components: probing asymmetric dynamical localization and measuring micromotion. To investigate localization phenomena, we implement the Bragg diffraction technique to launch the Bose-Einstein condensate (BEC) with varying recoil momentum, as outlined in [40]. By adjusting the frequency difference between the lattice beams and the on-time of the lattice beam, we achieve a good transfer of atoms to different momentum states. In our apparatus [41], a frequency difference of $15$ kHz results in $2$ recoil momenta, while a frequency difference of $30$ kHz provides $4$ recoil momenta, both have a fixed on-time of approximately $\sim 66.6$ $\mu$s. After the creation of the initial wave function of BEC with different velocities, we apply the usual kicked rotor pulse sequence to study the dynamical localization. In our QKR experiments, we maintain the stochastic parameter at $K=5$ and effective Planck constant at $\hbar_{\rm eff}=4.6$ to ensure that the corresponding classical dynamics remains in the chaotic regime [3].
For the motion of the lattice, we generate a frequency difference ranging from 00 to $75$ kHz between the lattice beams. This frequency range corresponds to a velocity range of 00 to $10$ recoil momentum. We ensure that the kick strength remains constant across higher frequency regimes. The advantage of employing a moving lattice, rather than a moving BEC, lies in the flexibility to assign precise arbitrary velocities to the optical lattice from the laboratory frame. It also does not create any residual atoms in zeroth momentum state like in Bragg diffraction. Leveraging this control over micromotion, we scan the lattice velocity in a frequency range of $-3$ to $3$ kHz with increments of $100$ Hz, enabling precise measurement of micromotion BEC by balancing the relative motion.
@@ IV.0.1 Case I: BEC moving in lab frame
In our experiment, BEC is launched with various recoil momenta, $p_{0}=n\hbar k$ ( $n\in\mathbb{Z}$ ), using Bragg diffraction. This is achieved by applying a pulse of length approximately $66.6$ $\mu$s to transfer all the population to the required momentum state, and appropriately adjusting the lattice power. Subsequently, free evolution period of approximately $66.6$ $\mu$s is allowed, corresponding to the Talbot time of the system. This ensures recreation of the initial wavefunction without any unintended phase accumulation [42]. Once the wavefunction with different velocities is created, it is subjected to periodic kicks to observe dynamical localization. The period of these kicks is set at $T=24.3$ $\mu$s, corresponding to a scaled Planck constant of $\hbar_{\rm eff}=4.6$. The stochastic parameter is $K=5$ corresponding to a classical phase space that is almost fully chaotic [37, 3]. These parameters remain consistent throughout the paper unless stated otherwise.
@@
Figure 1: Wavefunction profile for different launch velocity shown at $30$, $40$ and $50$ kicks. In all the cases, dynamical localization has set in. Initial velocity corresponds to momenta (a) $0\hbar k$, (b) $2\hbar k$, and (c) $4\hbar k$. In all cases, in top panel, symbols are obtained experimental data. In bottom panel, symbols are obtained from kicked rotor simulations with 50 kicks.
@@
Bose-Einstein Condensate is launched with initial momentum of $p_{0}=0\hbar k,2\hbar k$ and $4\hbar k$ corresponding, respectively to, zero, two and four recoil photons. The momentum distributions observed after $n=30,40$, and $50$ kicks are depicted in Fig. 1(a-c) respectively. In Fig. 1(a), standard symmetric dynamical localization pattern is observed for initial momentum $p_{0}=0$. Figure 1(b) illustrates an asymmetric dynamically localized momentum distribution peaked at $2\hbar k$ when BEC is launched with $p_{0}=2\hbar k$. Similarly, in Fig. 1(c), dynamical localization is asymmetric with peak positioned at $4\hbar k$, corresponding to a launch momentum of $p_{0}=4\hbar k$. Asymptotically, as $n\gg 1$, the system remembers the initial momentum $p_{0}$ since the maxima of the steady-state distribution occurs at $p=|p_{0}|$. As evident in Figs. 1(b,c), an asymmetry develops such that the localized momentum distribution in Fig. 1(b,c) has maxima at $p=-p_{0}$, when launched with initial momentum of $p_{0}$. In Fig. 1(b), $\langle p\rangle$ with reference to initial given momentum $p_{0}$, is moving in the direction of the launched velocity, while in Fig. 1(c), $\langle p\rangle$ is moving in the opposite direction of the launched velocity.
@@
The bottom panel in Fig. 1 shows the corresponding results obtained from QKR simulations. The simulation results confirm the emergence of dynamical localization and evidently it is asymmetric for the case when BEC is launched with $p_{0}=2\hbar k$ and $p_{0}=4\hbar k$. For $p_{0}=2\hbar k$, the more population lies in the direction of the launched momentum and for $p_{0}=4\hbar k$, it lies in opposite direction, matching with the experimental result. The experimental profile in the vicinity of the peak value, in an average sense, is slightly elevated compared to simulation results (in both Figs. 1-2) in due to presence residual thermal atoms.

### v3

@@ Abstract
We employ a Bose Einstein Condensate (BEC) based atom-optic kicked rotor to generate an asymmetrically localized momentum distribution that depends upon initial velocity of the BEC. Asymmetric features are shown to arise from the early-time dynamics induced by the broken parity symmetry and, asymptotically freeze as the dynamical localization stabilizes. The asymmetry in the momentum distribution critically depends upon the initial launch velocity and is sensitive to very small initial velocities (‘micromotion’) of the BEC. In this work, we also perform a precise measurement of the ‘micromotion’. By utilizing the technique of measuring the early-time asymmetry of momentum distribution, we report measurement of micromotion down to 230 $\pm$17 $\mu$m/s.
@@ I Introduction
In recent years, there has been much interest in studying quantum kicked rotor (QKR) with Bose-Einstein condensates (BEC) – both with and without tunable interactions [1, 2, 3, 4]. Further, the ease of imprinting the lattice phase on the BEC also leads to different types of effects such as on-resonance quantum ratchets [5, 6] and quantum boomerang effect [7, 8, 9]. BEC based QKR has also been utilized in coupled quantum kicked rotors [1] where two incommensurate optical lattices drive a quantum-to-classical transition by breaking dynamical localization (DL). Further, a discrete-time quantum walk has also been observed in a periodically kicked BEC-based system [10], and a variant with spin-1/2 degree of freedom has been proposed to realize and detect topological phase transitions [11, 12]. In this work, we utilize quantum kicked rotor experimentally realized with BEC in optical lattices for precision measurement of initial velocities towards applications in atom interferometry.
@@
Precision measurements with atom interferometers have opened up tremendous applications in quantum sensing [13, 14, 15]. Atom interferometers have been successfully utilized in gravimetry [16, 17, 15], rotation sensing [18, 19], magnetometers [20], and the determination of photon recoil [21]. In particular, in precision rotation sensing and gravimetry, nullifying the systematic shifts and measurement errors (arising from Coriolis effect [22]) necessitates precise knowledge of the initial velocity [18, 23, 16]. This initial velocity may arise from externally induced launch velocities or micromotion – small movement of the atomic cloud – during the turnoff of the trapping potential. In a rotation sensor, the phase shift caused due to this initial velocity adds to the Sagnac phase shift [22]. The Sagnac phase shift is in the range of tens of milliradians (mrad) for a velocity of 100 $\mu$m/s over an interferogram time of hundreds of milliseconds (ms). The velocity of micromotion can lie in the range of 100-1000 $\mu$m/s, which is much smaller compared to the velocity imparted due to two-photon recoil momentum ($\approx$ 1 cm/s). The measurement of such a small velocity is challenging because it requires huge time-of-flight in standard absorption imaging or very precise Bragg or Raman spectroscopy to measure the Doppler shift [23, 24, 25].
@@
In this work, we experimentally perform a precision measurement of ‘micromotion’ of the BEC. Firstly, we experimentally demonstrate asymmetric localization in momentum space by launching the BEC with varying initial velocity or by inducing the lattice motion. The accumulated phase difference imprinted on the launched wave packet or the lattice motion creates an asymmteric momentum distribution. Additionally, we illustrate that the asymmetric nature of the momentum distribution can be controlled by adjusting the launch velocity rather than altering the direction of the launch velocity. In contrast to previous studies [26] with cold atoms, where such asymmetry arises from pulse shape effects and observed after long time, here we demonstrate that such asymmetry arises much earlier in time and freezes due to dynamical localization. By utilizing the asymmetric behavior of the momentum distribution in the early-time dynamics, a method is presented to measure micromotion of the BEC [27, 28, 29] in QKR system. The QKR had already been utilized for measurement of gravity through survival resonances [30, 31] and this work adds another such example of precision measurement using BEC-based QKR.
II QKR in moving frame of reference
@@
The QKR is a fundamental model of quantum chaos [32] extensively explored for its chaotic dynamics and for demonstration of dynamical localization [33, 34]. The classical kicked rotor, when strongly kicked by an external field, can display chaotic dynamics accompanied by diffusive growth of mean energy with time. Over extended time scales, quantum interference effects inhibit classical diffusive dynamics, a phenomenon in momentum space analogous to Anderson localization [35]. Since the first realization of the QKR using cold atoms [36], it has spurred numerous experimental investigations to explore a variety of scenarios that manipulate localization [37, 38, 39]. Physically, the standard kicked rotor describes a particle subjected to periodic kicks imparted by the stationary optical lattice created by counter-propagating laser beams. However, in this work, we consider a kicked rotor system in which the optical lattice moves at a constant velocity in the laboratory frame. The Hamiltonian of QKR in a moving lattice is given by [40]
$H=\frac{\widehat{p}^{2}}{2}+K\cos(2k\widehat{x}-2\pi\alpha t)~{}\sum_{n=1}^{N}~{}\delta(t-nT),$ (1)
@@
where, $\widehat{p}$ and $\widehat{x}$ represent the momentum and position operators respectively. They obey canonical commutation relation $[\widehat{x},\widehat{p}]=i\hbar_{\rm eff}$, where the effective Planck constant $\hbar_{\rm eff}$ can be tuned in the experiment. Further, $K$ is the stochastic parameter, $T$ is the time period between consecutive kicks, $k$ is the wave vector, and $\alpha$ is the frequency difference between two lattice beams that make up the optical lattice. The lattice velocity arising from the frequency difference between the counter-propagating beams is $v=\frac{\lambda\alpha}{2}$, where $\lambda$ is the wavelength of the optical lattice. Throughout this work, parameters will be chosen so that the classical analogue of QKR displays chaos, with $K\geq 5$. This parameter choice ensures that the localization effects we observe are of quantum origin.
@@
$U=\exp{\left(-i\frac{p^{2}T}{2}\right)}\exp{\left(-iK\cos(2kx-2\pi\alpha t)\right)}.$ (2)
@@
and is consistent with the initial distribution of BEC. Here, $\sigma_{w}$ characterizes the width of the wavefunction in position space, while $p_{0}$ is the initial velocity, typically arising from launched velocity or micromotion. Generally, in our experiments, when BEC is launched with initial momentum $p_{0}$, the lattice velocity is stationary and vice versa. In general, either moving the lattice or the BEC in same direction are expected to induce similar effects in early time dynamics with exactly opposite asymmetric momentum distribution [41, 5]. The experimental protocol is illustrated in Fig. 1.
Figure 1: This experimental schematic shows the moving BEC (left side) with velocity $v$ and periodically kicked 1D optical lattice. This relative motion between BEC and lattice provides asymmetry in momentum distribution which we quantify by measuring $\langle p\rangle$. Same asymmetry with opposite nature arises when lattice is in motion, shown in right side. Micromotion is measured by scanning lattice velocity (see Fig. 5 for details). Further, dynamical localization sets in for larger number of kicks.
For numerical simulations, we use the standard split-operator method to evolve an initial state of the kicked rotor. This method consists of two primary components: the kick operator, which is diagonal in position space, and the free evolution operator, which is diagonal in momentum space. To simulate scenarios where the lattice is in motion in the lab frame, we imprint the phase just before each kick. Similarly, to model the situation in which BEC moves in the laboratory frame, a wavefunction is initialized incorporating the motion right from the outset.
@@
If the lattice is moved with constant velocity, then the kick potential (in the lab frame) $V(x)=K\cos(2kx-2\pi\alpha t)$ induces a path difference at each kick. Consequently, the phase difference between successive kicks is $\phi=2\pi\alpha T$ and this breaks the parity symmetry. The total phase difference accumulated after $n$ kicks is $\phi_{n}=2\pi\alpha T\times(n-1)$, assuming that the phase is initialized to zero for the first kick. The accumulated phase difference $\phi_{n}$ over short timescales induces an inhomogeneity in position space due to broken parity symmetry, leading to an asymmetric momentum distribution [5]. This asymmetry can be quantified through $\langle p(t=2T)\rangle$ immediately following $n=2$ kicks.
In the experiments, the lattice velocity $v$ is controlled by tuning $\alpha$ and subsequently measuring $\langle p(t=2T)\rangle$. Based on physical consideration and since phases are unique only upto $2\pi$, we posit that the average momentum after two kicks to have a form
$\langle p(t=2T)\rangle=c~{}\sin[4\pi(vT/\lambda)],$ (4)
where $c$ is a constant. Exactly the same expression holds good for moving BEC with opposite sign in average $\langle p\rangle$. Further, Eq. 4 implies that if $vT/\lambda=n/4$ (where $n$ is an integer), $\langle p\rangle=0$ implying an absence of asymmetry for specific choice of initial velocity and kick period. In the long time limit of $n\gg 1$, the initial asymmetry accumulated in the short-time limit eventually freezes due to the emergence of dynamical localization. Hence, the early-time dynamics dictates the long-term behavior and the onset of asymmetrical dynamical localization in the system.
To gauge micromotion accurately, the optical lattice is precisely moved with frequency difference of the order of $100$Hz, aligning it with the scale of micromotion. Upon achieving a velocity for the lattice that corresponds precisely to the micromotion, we observe that $\langle p(t=2T)\rangle$ equals to zero. This alignment establishes a direct correspondence between the velocity of the lattice and the velocity of micromotion.
@@ III Experimental setup for QKR
The QKR setup we have used for this work is similar to the one described in Ref. [39]. However, instead of cold atoms, we utilize a BEC of 87Rb every 8 seconds, through forced evaporative cooling. The atoms are initially prepared in the |$\ket{F=1,m_{F}=-1}$ state, with a BEC temperature of $80$ nK and a population of approximately $40,000$ atoms. This BEC serves as the initial wave function for our experimental investigations. The optical standing wave is produced using two independent laser beams, which are generated by a single laser passing through two separate acousto-optic modulators. The frequency difference between lattice beams and their switching can also be controlled. Switching time and laser power provide us control over scaled Planck constant $\hbar_{\rm eff}$ and stochastic parameter $K$.
This work comprises two main components: probing asymmetric dynamical localization and measuring micromotion. To investigate localization phenomena, we implement the Bragg diffraction technique to launch the Bose-Einstein condensate with varying recoil momentum, as outlined in Ref. [42]. By adjusting the frequency difference between the lattice beams and the on-time of the lattice beam, we achieve a good transfer of atoms to different momentum states. In our apparatus [43], a frequency difference of $15$ kHz results in $2$ photon recoil momenta, while a frequency difference of $30$ kHz provides $4$ photon recoil momenta to atoms, both have a fixed on-time of approximately $\sim 66.6$ $\mu$s. For a given frequency difference, optical lattice moves with half of the speed of diffracted wave-packets. After the creation of the initial wave function of BEC with different velocities, we apply the usual kicked rotor pulse sequence to study the dynamical localization. In our QKR experiments, we maintain the stochastic parameter at $K=5$ and effective Planck constant at $\hbar_{\rm eff}=4.6$ to ensure that the corresponding classical dynamics remains in the chaotic regime [3].
For the motion of the optical lattice, we generate a frequency difference ranging from 00 to $75$ kHz between the lattice beams. This frequency range corresponds to a velocity range:
$v=\frac{\alpha}{15kHz}.\frac{\hbar k}{M_{\text{Rb}}},$ (5)
where, $\hbar$k is single photon recoil momentum. We ensure that the kick strength remains constant across higher frequency regimes. The advantage of employing a moving lattice, rather than a moving BEC, lies in the flexibility to assign precise arbitrary velocities to the optical lattice from the laboratory frame. It also does not create any residual atoms in zeroth momentum state like in Bragg diffraction. Leveraging this control over micromotion, we scan the lattice velocity in a frequency range of $-3$ to $3$ kHz with increments of $100$ Hz, enabling precise measurement of micromotion BEC by balancing the relative motion.
@@ IV.0.1 Case I: BEC moving in lab frame
In our experiment, BEC is launched with various recoil velocity, $v_{n}=2n\hbar k/M_{Rb}$ ( $n\in\mathbb{Z}$ ), using Bragg diffraction, where $\hbar k/M_{Rb}$ is single photon recoil velocity. This is achieved by applying a pulse of length approximately $66.6$ $\mu$s to transfer all the population to the required momentum state, and appropriately adjusting the lattice power. Subsequently, free evolution period of approximately $66.6$ $\mu$s is allowed, corresponding to the Talbot time of the system [44]. This ensures recreation of the initial wavefunction without any unintended phase accumulation [45]. Once the wavefunction with different velocities is created, it is subjected to periodic kicks to observe dynamical localization. The period of these kicks is set at $T=24.3$ $\mu$s, corresponding to a scaled Planck constant of $\hbar_{\rm eff}=4.6$. The stochastic parameter is $K=5$ corresponding to a classical phase space that is almost fully chaotic [39, 3]. These parameters remain consistent throughout the paper unless stated otherwise.
@@
Figure 2: Wavefunction profile for different launch velocity shown at $30$, $40$ and $50$ kicks. In all the cases, dynamical localization has set in for initial velocity corresponds to momenta (a) $0\hbar k/M_{Rb}$, (b) $2\hbar k/M_{Rb}$, and (c) $4\hbar k/M_{Rb}$. In all cases, in left side, symbols are obtained experimental data. In right side, symbols are obtained from kicked rotor simulations with 50 kicks.
@@
Bose-Einstein Condensate is launched with initial velocity of $v_{0}=0,2\hbar k/M_{Rb}$ and $4\hbar k/M_{Rb}$. The momentum distributions observed after $30,40$, and $50$ kicks are depicted in Fig. 2(a-c) respectively in the left panel. In Fig. 2(a), standard symmetric dynamical localization pattern is observed for initial velocity $v_{0}=0$. Figure 2(b), illustrates an asymmetric dynamically localized momentum distribution when BEC is launched with $v_{0}=2\hbar k/M_{Rb}$ (peaked at $2\hbar k$, shifted to zero for better comparison). Similarly, in Fig. 2(c), dynamical localization is slightly asymmetric for a launch velocity of $v_{0}=4\hbar k/M_{Rb}$( peaked at $4\hbar k$, shifted to zero for better comparison). Asymptotically, as $n\gg 1$, the system remembers the initial velocity $v_{0}$ since the maxima of the steady-state distribution occurs at $p=|p_{0}|$. In Fig. 2(b), $\langle p\rangle$ with reference to initial given velocity $v_{0}$, is moving in the direction of the launched velocity (left direction), while in Fig. 2(c), small $\langle p\rangle$ is moving in the opposite direction of the launched velocity. As for the velocity of $v_{0}=4\hbar k/M_{Rb}$, $\langle p\rangle$ is small from the beginning, difficult to observe the asymmetry in localized state.
@@
The right panel in Fig. 2 shows the corresponding results obtained from QKR simulations. The simulation results confirm the emergence of dynamical localization and evidently it is asymmetric for the case when BEC is launched with $v_{0}=2\hbar k/M_{Rb}$ and $v_{0}=4\hbar k/M_{Rb}$. For $v_{0}=2\hbar k/M_{Rb}$, the more population lies in the direction of the launched velocity and for $v_{0}=4\hbar k/M_{Rb}$, it lies in opposite direction with little asymmetry, matching with the experimental result. This asymmetry feature in early-time dynamics will be discussed further in upcoming section. The experimental profile in the vicinity of the peak value, in an average sense, is slightly elevated compared to simulation results (in both Figs. 2-3) in due to presence residual thermal atoms.